In 2008, Ryan Newman won the Daytona 500 (mile) race with a speed of 152.672 mph. Find his time to the nearest thousandth.
The answer is 3.275 hr. But how do you determine the time?

Found 2 solutions by richwmiller, belladonnatink:
Answer by richwmiller(17219)

speed * time =distance
152.672*t=500
t=3.27499=3.275

Answer by belladonnatink(2)

152.672*t/152.672=500/152.672
Time=3.275